OpenStudy (anonymous):

As the marketing designer for Burger King, you are given a 4 in by 6 in logo and have been told to create an advertisement with sides six times larger. What is the area of your advertisement (in feet)?

OpenStudy (anonymous):

Start by finding the area of the smaller ad: 4*6=24 Multiply this by 6: 24*6=144 So the area is 144 inches squared. If there are 12 inches in a foot, then there are 12^2=144 square inches in a square foot. So the area of the new advert is one square foot.

OpenStudy (phi):

with sides six times larger that means multiply each side by 6 then change from inches to feet

OpenStudy (anonymous):

@emc do you understand?

OpenStudy (anonymous):

Yes, thanks!

OpenStudy (anonymous):

May I have a medal please :D

OpenStudy (anonymous):

ty very much